# SATURN: SAT-based Reinforcement Learning to Unleash Language Model Reasoning

### v3

@@ 1 Introduction
Recently, reinforcement learning (RL) has become a promising paradigm for unleashing the reasoning capability of large language models (LLMs), particularly in math, programming, and logical reasoning (e.g., OpenAI-o1 [openai_o1], DeepSeek-R1 [deepseek_r1], Kimi-k1.5 [kimi_k1.5]). During the RL training process, the design of RL tasks plays a critical role [openai_o3, openai_o3_web, s1]. A well-designed RL task should elicit LLMs’ reasoning capability, fostering behaviors such as hesitation, reflection, backtracking, summarization, and verification [cot, self_verification, stepback, s1, Verify_Step].
However, how to design RL tasks that can continuously enhance LLMs’ reasoning capability remains an open question. We think a well-designed RL task for reasoning should satisfy the following three criteria: ❶ Scalability. RL training requires large-scale data. RL tasks should support scalable data without human annotation or expensive LLMs’ synthesis. ❷ Verifiability. RL rewards must be unambiguously correct. The outputs of LLMs for the task should be easy to verify. ❸ Controllable Difficulty. Reasoning capability emerges progressively [Curriculum_Learning]. RL tasks should support the difficulty control to enable curriculum learning, allowing LLMs to gradually develop more complex reasoning skills [difficulty_controllable].
@@
ScaleQuest [question_synthesis]
GSM8K (Math) [gsm8k]
LiveCodeBench [lcb]
Game Werewolf [wolf1, wolf2]
LMRL Gym [lmrl]
SPAG [Self-playing-Adversarial]
Knights&Knaves [kk2]
@@
Table 1 shows the features of current mainstream RL tasks. None of them satisfy all three criteria. Existing RL tasks can be divided into two categories: (1) One category of RL tasks requires LLMs to solve math or programming problems, with rewards based on the correctness of the final answer or code [humaneval, aplhacode, mbpp, gsm8k]. However, these tasks rely on human annotation for ground-truth solutions or test cases, suffer from a lack of high-quality problems, and offer only coarse control over reasoning difficulty [aixcoder, aixcoderv2]. (2) Another category focuses on manually designed reasoning tasks [question_synthesis, wolf1, wolf2, kk2]. For instance, Logic-RL [Logic-RL] leverages natural language logic K&K puzzles to improve LLMs’ reasoning capability through RL. However, they also present limitations, such as hard to scale up due to reliance on sampling from LLMs [Self-playing-Adversarial, question_synthesis], hard to verify despite relying on LLMs for cross-validation [codedpo, question_synthesis, puzzbench], and hard to control difficulty [wolf1, wolf2].
@@
To this end, we propose Boolean Satisfiability (SAT) problem as the task for RL. Figure 1 shows an illustration of SAT problems and corresponding features. SAT satisfies all three desiderata we outlined earlier: ❶Scalability. SAT instances can be generated programmatically at scale without human annotation or LLM synthesis, allowing for virtually unlimited training data. ❷ Verifiability. SAT is a well-established NP-complete problem in theoretical computer science [SAT]. The correctness of a solution can be easily verified in linear time. But solving SAT problems requires complex reasoning. ❸ Controllable Difficulty. The difficulty of SAT instances can be precisely adjusted (e.g., number of variables, clauses), making it suitable for curriculum learning. What’s more, SAT serves as a universal substrate for limited forms of logical reasoning, as many problems in propositional logic, finite-domain first-order logic, and modal logic can be systematically reduced to SAT [logic1, logic2, logic3].
@@
We apply SATURN to DeepSeek-R1-Distill-Qwen-1.5B and 7B [deepseek_r1], obtaining SATURN-1.5B and SATURN-7B. Experiments show that SATURN effectively enhances LLMs’ reasoning capability in generalizable scenarios:
@@
  • The reasoning capability learned from SATURN transfers well to math and programming tasks, bringing average improvements of +4.9 and +1.8 on popular benchmarks such as AIME [aime], AMC [amc], MATH-500 [math500], GPQA Diamond [GPQA], and LiveCodebench [lcb] for two SATURN models.
@@ 2.3 Estimation of Task Difficulty
As a canonical NP-complete problem [SAT], SAT admits a polynomial-time reduction from any other NP problem [reduction]. SAT exhibits a known phase transition phenomenon: when the clause-to-variable ratio $\alpha_{c}=l/k$ approaches a critical threshold (typically near $4.26$ for 3-SAT), the probability of satisfiability drops sharply, and problem difficulty peaks. This phenomenon probably stems from replica symmetry breaking (RSB) [SAT-RSB]: near $\alpha_{c}$, the solution space fractures into disconnected clusters separated by energy barriers. Beyond $\alpha_{c}$, the space collapses, reducing complexity.
However, RSB theory is designed for heuristic SAT solvers. For humans or LLMs solving SAT problems through logical steps such as trial, verification, and reasoning, such solver-like phase transitions are hardly observable in human-like thinking processes. While any $n$-SAT ($n>3$) can be reduced to 3-SAT [reduction], they differ significantly for LLMs in terms of solution space size and token length.
Prior work [sat_iclr_1] on SAT tasks for LLMs typically categorized difficulty based on phase transition points. To systematically estimate task difficulty, we define an analytical estimator of the expected solution space size. Given a $(n,k,l)$ SAT instance, its difficulty for LLMs can be approximately estimated by:
@@ 2.4 Reinforcement Learning with GRPO
In this subsection, we introduce the single-stage RL training for given $(n,k,l)$-difficulty tasks. RL can further improve LLMs’ generalization by directly optimizing policy gradients over diverse reasoning trajectories [deepseek_r1]. Given the SAT tasks, we then train LLMs using the original sample-level GRPO to optimize the policy $\pi_{\theta}$ with KL divergence penalty. The GRPO objective function is defined as:
@@ 3.2 Experimental Setup
❶Building upon SAT_Construction tool and difficulty estimation, we release SATURN-2.6k, a curated benchmark designed to evaluate LLMs’ reasoning capability across varying complexity. SATURN-2.6k consists of 1,500 training instances and 160 test instances that share the same estimated difficulty level. To assess performance under increasing task complexity, SATURN-2.6k further includes 1,000 test instances from 10 unseen harder difficulty levels, with 100 instances per level. These levels are selected based on our difficulty estimation $D(n,k,l)$, enabling a systematic analysis of how LLM performance changes as problem difficulty increases. Additionally, custom datasets of desired difficulty can be constructed using our open-sourced SAT_Construction tool. ❷ For math and programming tasks, following DeepSeek-AI [deepseek_r1], we use AIME 24/25 [aime], AMC 22/23 [amc], MATH-500 [math500], GPQA Diamond [GPQA], and LiveCodeBench v4_v5 subset [lcb].
@@
We conduct evaluations against several 1.5B and 7B parameter reasoning models as the baselines, which include DeepSeek-R1-Distill-Qwen-1.5B & 7B [deepseek_r1], Still-3-1.5B-Preview [still3], s1.1-1.5B & 7B [s1], z1-7B [z1], OpenThinker-7B [openthoughts], and DeepScaleR-1.5B-Preview [deepscaler2025]. In addition, we include a supervised fine-tuning (SFT)-only baseline trained on the Math training dataset [math500], which provides step-by-step problem reasoning trajectories. We randomly select the most difficult Level-5 1,000 problems from training set for one epoch of SFT, following the same training template as DeepSeek-R1-Distill-Qwen. With the same dataset size, our setup enables a fair comparison between SFT and RL on SAT tasks.
@@
Following DeepSeek-AI [deepseek_r1], we use pass@k as the evaluation metric. Pass@k assesses the probability that at least one correct solution is generated within $k$ attempts. For SAT problems, we evaluate $\texttt{pass@k}\in\{1,3,5,7,10\}$ and sample 12 times per problem. For math and programming benchmarks, we use pass@1, following a context length of 32,768 and temperature = 0.6. More evaluation hyperparameters are provided in Appendix E. All experiments are conducted on NVIDIA 8×A100 (40GB) GPUs. Specific prompts are detailed in Appendix F.
@@ 3.4 RQ2: SATURN Demonstrates Strong Generalization to Math and Programming
❷SATURN outperforms SFT on broader benchmarks. Consistent with the observations in SFT Memorizes, RL Generalizes [SFT-RL], SFT improves performance on math-focused benchmarks (AIME, AMC, and Math500) that are similar to its supervised training domain. However, on LiveCodeBench, SFT drops from 16.4 to 14.6, exhibiting an alignment tax [alignment_tax], where specializing on a narrow domain compromises performance on other tasks. In contrast, SATURN improves performance across all benchmarks, with SATURN-1.5B reaching 17.4 on LiveCodeBench. Averaging across all benchmarks, SATURN-1.5B outperforms the SFT counterpart by 3.3, demonstrating that SATURN generalizes effectively.
@@ 3.5 RQ3: SATURN Serves as a Complement and Further Enhances LLM Reasoning
RQ3 studies the relationship between SATURN and existing RL tasks. Beyond the DeepSeek-R1-Distill-Qwen-7B, we introduce two additional models: Qwen2.5-7B-Instruct-1M [qwen2.5, qwen2.5-1m] following Logic-RL [Logic-RL] settings, and DeepScaleR-1.5B-Preview [deepscaler2025], which is further RL trained on 40k math and programming examples from DeepSeek-R1-Distill-Qwen-7B. We compare SATURN against several prior constructing RL task approaches, including Logic-RL [Logic-RL], SPGA [Self-playing-Adversarial], and ScaleQuest [question_synthesis], which represent strong baselines. Each approach is applied to different models for comparison. Results are summarized in Table 4.
@@ 3.6 RQ4: SATURN Enhances Self-verification in LLMs’ Reasoning Trajectories
RQ4 investigates whether SATURN affects LLMs’ reasoning behavior. On Qwen2.5-7B-Instruct-1M, we observe a gradual increase in response length during training, as illustrated in Figure 4, replicating the lengthening phenomenon reported in the R1 and Logic-RL [deepseek_r1, Logic-RL].
@@
Recent studies [Cognitive, Meta-Abilities] identify core behaviors shared by expert human reasoners and LLMs, such as verification and backtracking. These behaviors are domain-agnostic and provide fundamental reasoning patterns applicable to a wide range of tasks. In line with these findings, SATURN reinforces similar behaviors during SAT solving, leading to more structured reasoning trajectories. More reasoning trajectories are provided in Appendix J to illustrate how SATURN works. These results suggest that the self-verification patterns learned from SAT transfer well to math and programming tasks, improving reasoning robustness and reliability.
@@ 4.1 Limitations of Reasoning Capability Learned from SATURN
During curriculum learning, we observed that as the number of training iterations increases, the improvements in math and programming tasks tend to plateau, which is consistent with the findings in Logic-RL. Detailed evaluation results are provided in Table 5. This plateau may stem from several factors: ❶ Knowledge limitations. SATURN improves formal logical reasoning but does not provide domain-specific knowledge supervision. This limits its effectiveness in tasks requiring mathematical or algorithmic knowledge. ❷ Context window bottlenecks. SAT problems are NP-complete tasks, and the required reasoning length grows exponentially with increasing problem difficulty. This leads to bottlenecks in the model’s capability to handle increasingly complex tasks. ❸ Limited plasticity and forgetting. Model plasticity and catastrophic forgetting are known limitations that hinder further improvements with additional training stages [forgetting1, forgetting2].
@@ 5.1 Constructing Reasoning Tasks for Reinforcement Learning
Several works have explored constructing reasoning tasks to improve the reasoning capability of LLMs. Logic-RL [Logic-RL] and LMRL Gym [lmrl] train LLMs on natural language logic puzzles but lack scalability due to their limited puzzle set. ScaleQuest [question_synthesis], Entity-Deducing Game [20_Question], and K&K [kk2] propose automatic generation of constructing questions, but rely on LLM sampling or handcrafted templates, making large-scale generation costly and hard to verify. CodeDPO [codedpo] and PuzzBench [puzzbench] employ LLM-based verification, which may fail silently and cannot ensure correctness. Wolf Game [wolf1, wolf2] focus on multi-step logic reasoning but offer no control over task difficulty, limiting their support for curriculum learning. Overall, these tasks fall short in scalability, verifiability, or controllable difficulty. See Appendix K for detailed comparisons.
@@ 5.2 SAT-Based Evaluation of LLM Reasoning Capability
Recent studies have evaluated the reasoning capability of LLMs on SAT problems. Most of these works focus on analyzing model behavior around the SAT phase transition [sat_iclr_1, sat_iclr_solver, sat_blog], where problem hardness peaks. However, the phase transition theory is originally designed for heuristic SAT solvers and does not align well with the reflective and verification-based reasoning processes of humans or LLMs. These studies also lack a fine-grained scalable difficulty framework and typically divide difficulty based on the phase transition threshold. They are further limited to supervised fine-tuning and do not consider large reasoning models with long-CoT reasoning capability trained via RL. Our work addresses these limitations by building a progressive evaluation and curriculum learning pipeline, enabling precise difficulty control and the generalization of LLMs.
@@ Acknowledgements
This research is supported by the National Key R&D Program under Grant No. 2023YFB4503801, the National Natural Science Foundation of China under Grant No. 62192733, 62192730, 62192731, the Major Program (JD) of Hubei Province (No.2023BAA024).

### v4

@@ 1 Introduction
Recently, reinforcement learning (RL) has become a promising paradigm for unleashing the reasoning capability of large language models (LLMs), particularly in math, programming, and logical reasoning (e.g., OpenAI-o1 [21], DeepSeek-R1 [11], Kimi-k1.5 [37]). During the RL training process, the design of RL tasks plays a critical role [14, 32, 31]. A well-designed RL task should elicit LLMs’ reasoning capability, fostering behaviors such as hesitation, reflection, backtracking, summarization, and verification [43, 44, 53, 31, 28].
However, how to design RL tasks that can continuously enhance LLMs’ reasoning capability remains an open question. We think a well-designed RL task for reasoning should satisfy the following three criteria: ❶ Scalability. RL training requires large-scale data. RL tasks should support scalable data without human annotation or expensive LLMs’ synthesis. ❷ Verifiability. RL rewards must be unambiguously correct. The outputs of LLMs for the task should be easy to verify. ❸ Controllable Difficulty. Reasoning capability emerges progressively [42]. RL tasks should support the difficulty control to enable curriculum learning, allowing LLMs to gradually develop more complex reasoning skills [16].
@@
ScaleQuest [12]
GSM8K (Math) [9]
LiveCodeBench [22]
Game Werewolf [45, 48]
LMRL Gym [2]
SPAG [7]
Knights&Knaves [46]
@@
Table 1 shows the features of current mainstream RL tasks. None of them satisfy all three criteria. Existing RL tasks can be divided into two categories: (1) One category of RL tasks requires LLMs to solve math or programming problems, with rewards based on the correctness of the final answer or code [6, 27, 5, 9]. However, these tasks rely on human annotation for ground-truth solutions or test cases, suffer from a lack of high-quality problems, and offer only coarse control over reasoning difficulty [23, 26]. (2) Another category focuses on manually designed reasoning tasks [12, 45, 48, 46]. For instance, Logic-RL [47] leverages natural language logic K&K puzzles to improve LLMs’ reasoning capability through RL. However, they also present limitations, such as hard to scale up due to reliance on sampling from LLMs [7, 12], hard to verify despite relying on LLMs for cross-validation [51, 12, 41], and hard to control difficulty [45, 48].
@@
To this end, we propose Boolean Satisfiability (SAT) problem as the task for RL. Figure 1 shows an illustration of SAT problems and corresponding features. SAT satisfies all three desiderata we outlined earlier: ❶Scalability. SAT instances can be generated programmatically at scale without human annotation or LLM synthesis, allowing for virtually unlimited training data. ❷ Verifiability. SAT is a well-established NP-complete problem in theoretical computer science [10]. The correctness of a solution can be easily verified in linear time. But solving SAT problems requires complex reasoning. ❸ Controllable Difficulty. The difficulty of SAT instances can be precisely adjusted (e.g., number of variables, clauses), making it suitable for curriculum learning. What’s more, SAT serves as a universal substrate for limited forms of logical reasoning, as many problems in propositional logic, finite-domain first-order logic, and modal logic can be systematically reduced to SAT [17, 36, 24].
@@
We apply SATURN to DeepSeek-R1-Distill-Qwen-1.5B and 7B [11], obtaining SATURN-1.5B and SATURN-7B. Experiments show that SATURN effectively enhances LLMs’ reasoning capability in generalizable scenarios:
@@
  • The reasoning capability learned from SATURN transfers well to math and programming tasks, bringing average improvements of +4.9 and +1.8 on popular benchmarks such as AIME [3], AMC [4], MATH-500 [19], GPQA Diamond [35], and LiveCodebench [22] for two SATURN models.
@@ 2.3 Estimation of Task Difficulty
As a canonical NP-complete problem [10], SAT admits a polynomial-time reduction from any other NP problem [25]. SAT exhibits a known phase transition phenomenon: when the clause-to-variable ratio $\alpha_{c}=l/k$ approaches a critical threshold (typically near $4.26$ for 3-SAT), the probability of satisfiability drops sharply, and problem difficulty peaks. This phenomenon probably stems from replica symmetry breaking (RSB) [54]: near $\alpha_{c}$, the solution space fractures into disconnected clusters separated by energy barriers. Beyond $\alpha_{c}$, the space collapses, reducing complexity.
However, RSB theory is designed for heuristic SAT solvers. For humans or LLMs solving SAT problems through logical steps such as trial, verification, and reasoning, such solver-like phase transitions are hardly observable in human-like thinking processes. While any $n$-SAT ($n>3$) can be reduced to 3-SAT [25], they differ significantly for LLMs in terms of solution space size and token length.
Prior work [18] on SAT tasks for LLMs typically categorized difficulty based on phase transition points. To systematically estimate task difficulty, we define an analytical estimator of the expected solution space size. Given a $(n,k,l)$ SAT instance, its difficulty for LLMs can be approximately estimated by:
@@ 2.4 Reinforcement Learning with GRPO
In this subsection, we introduce the single-stage RL training for given $(n,k,l)$-difficulty tasks. RL can further improve LLMs’ generalization by directly optimizing policy gradients over diverse reasoning trajectories [11]. Given the SAT tasks, we then train LLMs using the original sample-level GRPO to optimize the policy $\pi_{\theta}$ with KL divergence penalty. The GRPO objective function is defined as:
@@ 3.2 Experimental Setup
❶Building upon SAT_Construction tool and difficulty estimation, we release SATURN-2.6k, a curated benchmark designed to evaluate LLMs’ reasoning capability across varying complexity. SATURN-2.6k consists of 1,500 training instances and 160 test instances that share the same estimated difficulty level. To assess performance under increasing task complexity, SATURN-2.6k further includes 1,000 test instances from 10 unseen harder difficulty levels, with 100 instances per level. These levels are selected based on our difficulty estimation $D(n,k,l)$, enabling a systematic analysis of how LLM performance changes as problem difficulty increases. Additionally, custom datasets of desired difficulty can be constructed using our open-sourced SAT_Construction tool. ❷ For math and programming tasks, following DeepSeek-AI [11], we use AIME 24/25 [3], AMC 22/23 [4], MATH-500 [19], GPQA Diamond [35], and LiveCodeBench v4_v5 subset [22].
@@
We conduct evaluations against several 1.5B and 7B parameter reasoning models as the baselines, which include DeepSeek-R1-Distill-Qwen-1.5B & 7B [11], Still-3-1.5B-Preview [40], s1.1-1.5B & 7B [31], z1-7B [50], OpenThinker-7B [38], and DeepScaleR-1.5B-Preview [29]. In addition, we include a supervised fine-tuning (SFT)-only baseline trained on the Math training dataset [19], which provides step-by-step problem reasoning trajectories. We randomly select the most difficult Level-5 1,000 problems from training set for one epoch of SFT, following the same training template as DeepSeek-R1-Distill-Qwen. With the same dataset size, our setup enables a fair comparison between SFT and RL on SAT tasks.
@@
Following DeepSeek-AI [11], we use pass@k as the evaluation metric. Pass@k assesses the probability that at least one correct solution is generated within $k$ attempts. For SAT problems, we evaluate $\texttt{pass@k}\in\{1,3,5,7,10\}$ and sample 12 times per problem. For math and programming benchmarks, we use pass@1, following a context length of 32,768 and temperature = 0.6. More evaluation hyperparameters are provided in Appendix E. All experiments are conducted on NVIDIA 8×A100 (40GB) GPUs. Specific prompts are detailed in Appendix F.
@@ 3.4 RQ2: SATURN Demonstrates Strong Generalization to Math and Programming
❷SATURN outperforms SFT on broader benchmarks. Consistent with the observations in SFT Memorizes, RL Generalizes [8], SFT improves performance on math-focused benchmarks (AIME, AMC, and Math500) that are similar to its supervised training domain. However, on LiveCodeBench, SFT drops from 16.4 to 14.6, exhibiting an alignment tax [33], where specializing on a narrow domain compromises performance on other tasks. In contrast, SATURN improves performance across all benchmarks, with SATURN-1.5B reaching 17.4 on LiveCodeBench. Averaging across all benchmarks, SATURN-1.5B outperforms the SFT counterpart by 3.3, demonstrating that SATURN generalizes effectively.
@@ 3.5 RQ3: SATURN Serves as a Complement and Further Enhances LLM Reasoning
RQ3 studies the relationship between SATURN and existing RL tasks. Beyond the DeepSeek-R1-Distill-Qwen-7B, we introduce two additional models: Qwen2.5-7B-Instruct-1M [49, 39] following Logic-RL [47] settings, and DeepScaleR-1.5B-Preview [29], which is further RL trained on 40k math and programming examples from DeepSeek-R1-Distill-Qwen-7B. We compare SATURN against several prior constructing RL task approaches, including Logic-RL [47], SPGA [7], and ScaleQuest [12], which represent strong baselines. Each approach is applied to different models for comparison. Results are summarized in Table 4.
@@ 3.6 RQ4: SATURN Enhances Self-verification in LLMs’ Reasoning Trajectories
RQ4 investigates whether SATURN affects LLMs’ reasoning behavior. On Qwen2.5-7B-Instruct-1M, we observe a gradual increase in response length during training, as illustrated in Figure 4, replicating the lengthening phenomenon reported in the R1 and Logic-RL [11, 47].
@@
Recent studies [15, 20] identify core behaviors shared by expert human reasoners and LLMs, such as verification and backtracking. These behaviors are domain-agnostic and provide fundamental reasoning patterns applicable to a wide range of tasks. In line with these findings, SATURN reinforces similar behaviors during SAT solving, leading to more structured reasoning trajectories. More reasoning trajectories are provided in Appendix J to illustrate how SATURN works. These results suggest that the self-verification patterns learned from SAT transfer well to math and programming tasks, improving reasoning robustness and reliability.
@@ 4.1 Limitations of Reasoning Capability Learned from SATURN
During curriculum learning, we observed that as the number of training iterations increases, the improvements in math and programming tasks tend to plateau, which is consistent with the findings in Logic-RL. Detailed evaluation results are provided in Table 5. This plateau may stem from several factors: ❶ Knowledge limitations. SATURN improves formal logical reasoning but does not provide domain-specific knowledge supervision. This limits its effectiveness in tasks requiring mathematical or algorithmic knowledge. ❷ Context window bottlenecks. SAT problems are NP-complete tasks, and the required reasoning length grows exponentially with increasing problem difficulty. This leads to bottlenecks in the model’s capability to handle increasingly complex tasks. ❸ Limited plasticity and forgetting. Model plasticity and catastrophic forgetting are known limitations that hinder further improvements with additional training stages [1, 13].
@@ 5.1 Constructing Reasoning Tasks for Reinforcement Learning
Several works have explored constructing reasoning tasks to improve the reasoning capability of LLMs. Logic-RL [47] and LMRL Gym [2] train LLMs on natural language logic puzzles but lack scalability due to their limited puzzle set. ScaleQuest [12], Entity-Deducing Game [52], and K&K [46] propose automatic generation of constructing questions, but rely on LLM sampling or handcrafted templates, making large-scale generation costly and hard to verify. CodeDPO [51] and PuzzBench [41] employ LLM-based verification, which may fail silently and cannot ensure correctness. Wolf Game [45, 48] focus on multi-step logic reasoning but offer no control over task difficulty, limiting their support for curriculum learning. Overall, these tasks fall short in scalability, verifiability, or controllable difficulty. See Appendix K for detailed comparisons.
@@ 5.2 SAT-Based Evaluation of LLM Reasoning Capability
Recent studies have evaluated the reasoning capability of LLMs on SAT problems. Most of these works focus on analyzing model behavior around the SAT phase transition [18, 34, 30], where problem hardness peaks. However, the phase transition theory is originally designed for heuristic SAT solvers and does not align well with the reflective and verification-based reasoning processes of humans or LLMs. These studies also lack a fine-grained scalable difficulty framework and typically divide difficulty based on the phase transition threshold. They are further limited to supervised fine-tuning and do not consider large reasoning models with long-CoT reasoning capability trained via RL. Our work addresses these limitations by building a progressive evaluation and curriculum learning pipeline, enabling precise difficulty control and the generalization of LLMs.
@@ Acknowledgements
This research is supported by the National Key R&D Program under Grant No. 2023YFB4503801, the National Natural Science Foundation of China under Grant No. 62192733, 62192730, 62192731, the Major Program (JD) of Hubei Province (No.2023BAA024), and the Beijing Natural Science Foundation under Grant No. 4264107.